OpenStudy (anonymous):

x + 2y = 50 -2x + 3y = 40 5x + 5y +5z = 300

OpenStudy (anonymous):

(-10, -20, -30) (10, 20 ,30) (10,20, -30) (10, -20, 30) What would be the answer?

OpenStudy (anonymous):

@Mertsj could you please help me?

OpenStudy (anonymous):

@Hero would it be a problem for you?

OpenStudy (anonymous):

@ganeshie8 ??

OpenStudy (anonymous):

im begging

hero (hero):

Divide the both sides of 5x + 5y + 5z = 300 by 5 to get: x + y + z = 60 So now you have: x + 2y = 50 -2x + 3y = 40 x + y + z = 60 Pair the 1st and second equations together: x + 2y = 50 -2x + 3y = 40 x + 2y = 50 by 2 to get: 2x + 4y = 100 -2x + 3y = 40 Then combine them together to get: 7y = 140 y = 20 Since y = 20, use that to find x: x + 2y = 50 x + 2(20) = 50 x + 40 = 50 x = 50 - 40 x = 10 Since x =10, y = 20, find z: x + y + z = 60 10 + 20 + z = 60 30 + z = 60 z = 60 - 30 z = 30 So (x,y,z) = (10, 20, 30)